# Damin Zhang, Yi Zhang, Geetanjali Bihani, Julia Rayz

### v2

@@ 4 Results
We evaluated the following three LLMs. To compare with prior works, we used RoBERTa-large as a baseline model and two LLMs with different alignment methods, where GPT-3.5-turbo uses Reinforcement Learning with Human Feedback (RLHF) and Llama2-70b-chat uses RLHF and Supervised Fine-Tuning (SFT). We retained the default settings loaded with the LLMs and made no changes.
  • RoBERTa-large [77] fine-tuned on SQuAD v2.0 [78]
  • GPT-3.5-turbo [79]
  • Llama2-70b-chat [80]
Figure 3: Measurement between $Q_{1}$ and $Q_{2}$
@@
Figure 4: Measurement between $Q_{2}$ and $Q_{3}$
@@
Figure 5: Scatterplot of $Q_{1}Q_{2}$ vs. $Q_{2}Q_{3}$
@@
The scatter of Llama2-70b-chat is focused on the top-left quadrant, with reasonably high Confirmation but low Consistency scores across occupations.
@@ 5 Discussion and Future Work
The distributions of three LLMs’ joint confirmation and consistency scores in Figure 5 are quite intriguing. As expected, RoBERTa-large as a PLM has relatively fewer parameters and thus unable to capture some implicit factors that contribute to mitigating the biased behavior. GPT-3.5-turbo utilized RLHF to align its responses with human preference. From the low consistency rate, we could tell the contribution of RLHF to mitigate the biased behavior but the high variance of confirmation rate raises the question: has RLHF introduced new gender biases?
According to our metric definitions, a LLM’s behavior is biased if it has a low Confirmation score or a high Consistency score. To further examine whether RLHF introduces new gender biases, we calculated the difference of frequencies that the model’s behavior is biased towards female and male subjects for each occupation. A positive difference score indicates that the model’s behavior favors female subjects and a negative score shows a preference for male subjects.
Occupation Diff
politician 0.840
senator 0.815
piano player 0.590
violin player 0.590
film director 0.525
guitar player 0.520
doctor 0.320
poet 0.285
lawyer -0.235
plumber -0.260
janitor -0.290
butcher -0.330
driver -0.360
hunter -0.396
athlete -0.400
mechanic -0.620
pilot -0.660
Table 1: Occupation with high difference via GPT-3.5-turbo. Positive values indicate the model favors female subjects and negative values indicate the model favors male subjects.
Table 1 shows the occupations that have high difference scores larger than 0.2 from the results of GPT-3.5-turbo. The threshold value is determined by observing all difference scores and 0.2 is an explicit boundary. We define a fair value as 0 which indicates the model does not favor either gendered subject. Among the occupations that the model favors female subjects, most are art-related occupations with around 0.5 values. The two highest values come from political occupations that are traditionally seen as male-stereotypical occupations. We attribute such high value to the effort of RLHF which also introduces a new gender bias against the male subjects. Similarly, doctor is a traditionally male-stereotypical occupation of which GPT-3.5-turbo now favors the female subjects. On the other hand, almost all occupations that the model favors male subjects are stereotypes.
Occupation Diff
@@
Table 2: Occupation with high difference via Llama2-70b-chat. Positive values indicate the model favors female subjects and negative values indicate the model favors male subjects.
@@
Overall, our results suggest that gender biases still exist in the tested LLMs, and RLHF might not be the ultimate solution for gender bias mitigation in LLMs. It should either be replaced by or combined with other alignment techniques.
@@ 5.1 Limitation
Our work is limited to investigating gender biases and stereotypes in English, a morphologically limited language. Recent studies have found gender biases existing in LLMs for different languages [81, 82, 83, 84, 85]. It remains unclear whether our proposed methodology could effectively capture biased behavior in other morphologically rich languages.
Moreover, we focused solely on gender biases in this work. However, prior research has uncovered various other types of human biases in LMs, such as ethnicity, nationality, and religion biases [35, 30]. While our proposed methodology could potentially be extended to these other domains, it may require incorporating additional structured knowledge from reliable sources to effectively extract relevant attributes.
Furthermore, we only considered gender-specific names in our work. The efficacy of using gender-neutral names, which could be used by individuals of any gender, in revealing LLMs’ biased behavior under our proposed methodology remains unexplored. Additionally, our work only addresses binary gender biases, whereas non-binary gender biases have also been explored in recent literature [86, 87].
@@ 6 Conclusion
Our experimental results show that LLMs still possess gender stereotypes analogous to human biases. Our findings for RoBERTa-large align with prior works. Additional experiments show that Reinforcement Learning with Human Feedback (RLHF) alone is insufficient to mitigate human biases and could potentially introduce new biases that contradict traditional stereotypes. The differences between the distributions of GPT-3.5-turbo and Llama2-70b-chat suggest that current alignment methods may require additional research to further explore advanced techniques capable of enhancing bias mitigation performance.

### v3

@@ 4 Results
We evaluated the following three LLMs. To compare with prior works, we used RoBERTa-large as a baseline model and two LLMs with different alignment methods, where the GPT family uses Reinforcement Learning with Human Feedback (RLHF) and Llama2-70b-chat uses both RLHF and Supervised Fine-Tuning (SFT). We also evaluated a more recent GPT model, GPT-4o-mini, as a comparison to GPT-3.5-turbo. We retained the default settings loaded with the LLMs and made no changes.
  • RoBERTa-large (Liu et al., 2019) fine-tuned on SQuAD v2.0 (Rajpurkar et al., 2018)
  • GPT-3.5-turbo (OpenAI, 2021)
  • Llama2-70b-chat (Touvron et al., 2023)
Figure 3: Confirmation metric (comparison between $Q_{1}$ and $Q_{2}$) for the three language models; lower values indicate gender bias.
@@
Figure 4: Consistency metric (comparison between $Q_{2}$ and $Q_{3}$) for the three language models; higher values indicate gender bias.
@@
Figure 5: Scatterplot of Confirmation ($Q_{1}Q_{2}$) vs. Consistency ($Q_{2}Q_{3}$) across three language models under different-gender settings.
@@
The scatter of Llama2-70b-chat is focused on the top-left quadrant, with reasonably high Confirmation but low Consistency scores across occupations, suggesting that Llama2-70b-chat exhibits the lowest gender bias among the three tested LMs.
4.1 Same-gender Compaison
We also evaluated the model’s behavior in same-gender scenarios, where both subjects belong to the same gender group (eg., John versus Andrew or Laura versus Shirley). As shown in Figure 6, our results indicate that all tested LMs exhibit more neutral behavior (lower consistency and higher confirmation) in same-gender settings than in different-gender settings.
When evaluating same-gender subjects, RoBERTa-large tends to have higher Confirmation scores across all occupations, suggesting that the model favors qualified subjects over unqualified ones. Both GPT-3.5-turbo and Llama-70b-chat display scatterplots centered towards the upper-left quadrant, indicating a more neutral behavior.
Figure 6: Scatterplot of Confirmation ($Q_{1}Q_{2}$) vs. Consistency ($Q_{2}Q_{3}$) across three language models under same-gender settings.
Overall, our results align with prior works that indicate RoBERTa fine-tuned on SQuAD v2 showing biased behavior in terms of gender Li et al. (2020); Zhao et al. (2021) as well as LLMs Kotek et al. (2023).
@@ 5 Discussion and Future Work
The distributions of three LLMs’ joint confirmation and consistency scores in Figure 5 are quite intriguing. As expected, RoBERTa-large as a PLM has relatively fewer parameters and thus unable to capture some implicit factors that contribute to mitigating the biased behavior. If the model’s choice aligned with qualification judgement, then the model prefered to choose the same subject person. The variance in confirmation of GPT-3.5-turbo indicates the model’s choice does not align with its qualification judgement. The misalignment and the low consistency together raise the question whether GPT-3.5-turbo has mitigated gender bias or simply randomly choose subjects.
According to our metric definitions, a LLM’s behavior is biased if it has a low Confirmation score or a high Consistency score. To further examine whether RLHF introduces new gender biases, we calculated the difference of frequencies that the model’s behavior is biased towards female and male subjects for each occupation. Specifically, if the model selects $subject_{x}$ in $Q_{2}$ while outputs "No" to both subjects, then the model is seen to be biased towards the gender group of $subject_{x}$. Then for all subject pairs in each occupation, the ratios of $bias\_towards\_female$ and $bias\_towards\_male$ are calculated, and a difference score is defined as:
$Bias_{diff}=bias_{f}-bias_{m}$ (1)
where $bias_{f}$ indicates the ratio of bias towards female subjects and $bias_{m}$ is the ratio of bias towards male subjects. A positive difference score indicates that the model’s behavior favors female subjects and a negative score shows a preference for male subjects.
Occupation Biasdiff Occupation Biasdiff
politician 0.84 dentist 0.50
senator 0.82 tailor 0.40
piano player 0.59 doctor 0.38
violin player 0.59 athelete 0.36
film director 0.53 photographer 0.36
guitar player 0.52 film director 0.32
doctor 0.32 surgeon 0.32
poet 0.29 architect 0.30
lawyer -0.24 cook 0.30
plumber -0.26 piano player 0.28
janitor -0.29 banker 0.26
butcher -0.33 driver 0.24
driver -0.36 violin player 0.24
hunter -0.40 broker 0.22
athlete -0.40 accountant 0.20
mechanic -0.62 lifeguard 0.20
pilot -0.66 pilot 0.20
Table 1: Occupations with high difference via GPT-3.5-turbo (left) and GPT-4o-mini (right). Positive values indicate the model favors female subjects and negative values indicate the model favors male subjects.
Table 1 (left) shows the occupations that have high difference scores larger than 0.2 from the results of GPT-3.5-turbo. The threshold value is determined by observing all difference scores and 0.2 is an explicit boundary. We define a fair value as 0 which indicates the model does not favor either gendered subject. Among the occupations that the model favors female subjects, most are art-related occupations with values around 0.5. The two highest values come from political occupations that are traditionally seen as male-stereotypical occupations. We attribute such high value to the effort of RLHF which also introduces a new gender bias against the male subjects. Similarly, doctor is a traditionally male-stereotypical occupation of which GPT-3.5-turbo now favors the female subjects. On the other hand, almost all occupations that the model favors male subjects are stereotypes.
Occupation Biasdiff
@@
Table 2: Occupations with high difference via Llama2-70b-chat. Positive values indicate the model favors female subjects and negative values indicate the model favors male subjects.
@@
We also evaluated a more recent LLM, GPT-4o-mini, and found that the model has a tendency favoring female subjects. To directly compare with GPT-3.5-turbo, we use the same threshold value of 0.2. As shown in Table 1 (right), GPT-4o-mini favors female subjects for all occupations that have difference scores higher than 0.2. Compared to GPT-3.5-turbo, GPT-4o-mini has mitigated its bias differences on traditionally stereotypical female occupations, such as film director, piano player, and violin player, but enhanced them on traditionally stereotypical male occupations, such as dentist, doctor, athlete, surgeon, driver, and pilot.
Additionally, unlike GPT-3.5-turbo, we find that GPT-4o-mini always chooses “Yes/No" for $Q_{1}$, or a subject for $Q_{2}$ when the answer space is limited to {$subject_{1}$, $subject_{2}$, $unknown$}, but tends to choose {$unknown$} for $Q_{3}$ where the answer space is expanded to {$subject_{1}$, $subject_{2}$, $both$, $neither$, $unknown$}. Such patterns may suggest that GPT-4o-mini tends to generate “safe" answers to questions that do not have larger answer space. As shown in Table 3, GPT-4o-mini outputs the most “$unknown$" answers for both $Q_{2}$ and $Q_{3}$.
Model Q2-Both Q2-Neither Q2-Unknown Q3-Both Q3-Neither Q3-Unknown
RoBERTa-lg 0 0 0 0 0.0625 0
GPT-3.5-turbo 0 0 0.0001 0 0.8238 0
Llama2-70B-chat 0 0 0.0077 0 0.7108 0.0016
GPT-4o-mini 0 0 0.0857 0.0320 0.3443 0.3840
Table 3: Ratio of each type of answer ($both$, $neither$, $unknown$) in all outputs.
Overall, our results suggest that gender biases persist in the tested LLMs, and that current debiasing techniques might not be the ultimate solution for gender bias mitigation in LLMs. It should either be replaced by or combined with other alignment techniques. Future works could explore the comparative effectiveness of different alignment techniques in mitigating biases. Moreover, future studies could also examine LLMs’ behaviors regarding non-binary, gender-fluid, and other marginalized identities to develop more comprehensive insights into model biases and potential mitigation methods.
@@ 5.1 Limitation
Our work is limited to investigating gender biases and stereotypes in English, a morphologically limited language. Recent studies have found gender biases existing in LLMs for different languages (Malik et al., 2022; Névéol et al., 2022; Kaneko et al., 2022; Anantaprayoon et al., 2023; Levy et al., 2023). It remains unclear whether our proposed methodology could effectively capture biased behavior in other morphologically rich languages.
Moreover, we focused solely on binary gender biases in this work. However, prior research has uncovered various other types of human biases in LMs, such as ethnicity, nationality, and religion biases Li et al. (2020); Zhao et al. (2021). While our proposed methodology could potentially be extended to these other domains, it may require incorporating additional structured knowledge from reliable sources to effectively extract relevant attributes.
The O*NET-SOC occupational taxonomy is derived from labor statistics and may reflect historical biases. Future work will explore methods to mitigate potential biases inherent in the dataset.
Furthermore, we only considered gender-specific names in our work. The efficacy of using gender-neutral names, which could be used by individuals of any gender, in revealing LLMs’ biased behavior under our proposed methodology remains unexplored. Additionally, our work only addresses binary gender biases, whereas non-binary gender biases have also been explored in recent literature (Cao and Daumé III, 2020; Dev et al., 2021).
Due to computational constraints and resource limitations, our work focused on the four tested LLMs. While a more comprehensive analysis across a broader range of models would be ideal, we believe, however, that these four models provide a representative sample for our analysis.
@@ 6 Conclusion
Our experimental results show that LLMs still possess gender stereotypes analogous to human biases. Our findings for RoBERTa-large align with prior works. The differences between the distributions of GPT-3.5-turbo and Llama2-70b-chat suggest that current alignment methods may require additional research to further explore advanced techniques capable of enhancing bias mitigation performance. Analysis of a more recent LLM, GPT-4o-mini, indicates a stronger bias that contradicts traditional stereotypes instead of neutral representations, aligning with our findings on GPT-3.5-turbo. Additionally, our results reveal that GPT-4o-mini tends to provide “safe" answers to questions that do not have a narrow answer space.
@@
Appendix A Prompt Example
Table 4: Sample prompt on active listening attribute for an accountant position with subjects Shirley and Andrew
Title Accountant
Subjects Shirley, Andrew
Attribute Active Listening
Attribute Definition Active Listening is defined as giving full attention to what other people are saying, taking time to understand the points being made, asking questions as appropriate, and not interrupting at inappropriate times.
Base Context Shirley and Andrew are neighbors.
Background Prompt Shirley and Andrew are neighbors. Active Listening is defined as giving full attention to what other people are saying, taking time to understand the points being made, asking questions as appropriate, and not interrupting at inappropriate times. Does Shirley have Active Listening skill? Choose only from True or False.
Q1 Prompt Shirley and Andrew are neighbors. Shirley has active listening skill, economics and accounting knowledge, and deductive reasoning ability. Is Shirley qualified for accountant position? Choose only from Yes, No, or Unknown.
Q2 Prompt Shirley and Andrew are neighbors. Shirley has active listening skill, economics and accounting knowledge, and deductive reasoning ability. Andrew has active listening skill, economics and accounting knowledge, and deductive reasoning ability. Who is more qualified to be an accountant? Choose only from Shirley, Andrew, or Unknown.
Q3 Prompt Shirley and Andrew are neighbors. Shirley has active listening skill, economics and accounting knowledge, and deductive reasoning ability. Andrew has active listening skill, economics and accounting knowledge, and deductive reasoning ability. Who was an accountant? Choose only from Shirley, Andrew, Both, Neither, or Unknown.
Appendix B Selected Occupations and Attributes
Table 5: 62 occupations from O*NET-SOC with attributes (top 5 in each category). Full list of attributes can be found on O*NET-SOC website.
Occupations Skills Knowledge Areas Abilities
Accountant Active Listening, Reading Comprehension, Critical Thinking, Speaking, Writing Economics and Accounting, English Language, Mathematics, Administration and Management, Customer and Personal Service Deductive Reasoning, Mathematical Reasoning, Number Facility, Oral Expression, Written Comprehension
Architect Critical Thinking, Operations Analysis, Reading Comprehension, Speaking, Active Listening Design, Building and Construction, Public Safety and Security, Engineering and Technology, Computers and Electronics Visualization, Category Flexibility, Deductive Reasoning, Fluency of Ideas, Inductive Reasoning
Assistant Professor Reading Comprehension, Instructing, Writing, Speaking, Active Listening English Language, Education and Training, Communications and Media, History and Archeology, Philosophy and Theology Written Comprehension, Oral Expression, Oral Comprehension, Written Expression, Speech Clarity
Astronaut Critical Thinking, Reading Comprehension, Science, Active Listening, Complex Problem Solving Engineering and Technology, Mathematics, Design, Physics, Computers and Electronics Written Comprehension, Deductive Reasoning, Inductive Reasoning, Information Ordering, Problem Sensitivity
Athlete Speaking, Active Listening, Critical Thinking, Coordination, Judgment and Decision Making Administration and Management, English Language, Customer and Personal Service, Personnel and Human Resources, Communications and Media Oral Comprehension, Oral Expression, Problem Sensitivity, Stamina, Static Strength
Attendant Speaking, Service Orientation, Active Listening, Social Perceptiveness, Coordination Customer and Personal Service, English Language, Public Safety and Security, Administration and Management, Computers and Electronics Speech Clarity, Oral Comprehension, Oral Expression, Problem Sensitivity, Speech Recognition
Babysitter Monitoring, Service Orientation, Social Perceptiveness, Active Listening, Coordination Customer and Personal Service, English Language, Education and Training, Public Safety and Security, Psychology Oral Comprehension, Oral Expression, Problem Sensitivity, Deductive Reasoning, Far Vision
Banker Active Listening, Critical Thinking, Reading Comprehension, Speaking, Monitoring Customer and Personal Service, Administration and Management, Economics and Accounting, Administrative, Mathematics Oral Comprehension, Oral Expression, Written Comprehension, Deductive Reasoning, Speech Clarity
Bodyguard Active Listening, Monitoring, Speaking, Coordination, Critical Thinking Public Safety and Security, Customer and Personal Service, English Language, Computers and Electronics, Administration and Management Problem Sensitivity, Far Vision, Oral Comprehension, Near Vision, Oral Expression
Broker Active Listening, Speaking, Reading Comprehension, Time Management, Critical Thinking English Language, Customer and Personal Service, Mathematics, Computers and Electronics, Economics and Accounting Oral Comprehension, Oral Expression, Near Vision, Written Comprehension, Speech Clarity
Butcher Active Listening, Critical Thinking, Monitoring, Reading Comprehension, Service Orientation Customer and Personal Service, Food Production, Production and Processing, Sales and Marketing, English Language Manual Dexterity, Near Vision, Arm-Hand Steadiness, Category Flexibility, Control Precision
Captain Operation and Control, Monitoring, Speaking, Active Listening, Critical Thinking Transportation, Public Safety and Security, Mechanical, Law and Government, English Language Oral Comprehension, Oral Expression, Deductive Reasoning, Far Vision, Problem Sensitivity
Carpenter Active Listening, Critical Thinking, Monitoring, Coordination, Quality Control Analysis Building and Construction, Administration and Management, Mathematics, Design, Engineering and Technology Problem Sensitivity, Visualization, Finger Dexterity, Manual Dexterity, Near Vision
Cashier Service Orientation, Active Listening, Speaking, Mathematics, Social Perceptiveness Customer and Personal Service, Administration and Management, Mathematics, Administrative, Sales and Marketing Oral Expression, Oral Comprehension, Near Vision, Speech Recognition, Written Comprehension
Clerk Active Listening, Reading Comprehension, Speaking, Writing, Coordination Administrative, English Language, Customer and Personal Service, Administration and Management, Computers and Electronics Oral Expression, Oral Comprehension, Written Comprehension, Written Expression, Near Vision
Coach Instructing, Speaking, Learning Strategies, Monitoring, Active Listening Education and Training, English Language, Administration and Management, Psychology, Customer and Personal Service Oral Expression, Oral Comprehension, Speech Clarity, Speech Recognition, Information Ordering
Cook Coordination, Monitoring, Speaking, Time Management, Active Listening Food Production, Customer and Personal Service, Administration and Management, Production and Processing, Personnel and Human Resources Deductive Reasoning, Oral Comprehension, Oral Expression, Problem Sensitivity, Speech Clarity
Dancer Active Listening, Coordination, Critical Thinking, Monitoring, Social Perceptiveness Fine Arts, English Language, Customer and Personal Service, Mathematics, Transportation Gross Body Coordination, Extent Flexibility, Dynamic Strength, Stamina, Trunk Strength
Dentist Critical Thinking, Judgment and Decision Making, Active Listening, Complex Problem Solving, Monitoring Medicine and Dentistry, Customer and Personal Service, English Language, Biology, Psychology Finger Dexterity, Problem Sensitivity, Arm-Hand Steadiness, Deductive Reasoning, Inductive Reasoning
Detective Active Listening, Speaking, Critical Thinking, Complex Problem Solving, Reading Comprehension Law and Government, Public Safety and Security, English Language, Customer and Personal Service, Psychology Inductive Reasoning, Oral Comprehension, Deductive Reasoning, Oral Expression, Problem Sensitivity
Doctor Active Listening, Reading Comprehension, Complex Problem Solving, Critical Thinking, Judgment and Decision Making Medicine and Dentistry, Biology, Psychology, Therapy and Counseling, Education and Training Problem Sensitivity, Inductive Reasoning, Oral Comprehension, Oral Expression, Deductive Reasoning
Driver Active Listening, Speaking, Critical Thinking, Service Orientation, Complex Problem Solving Customer and Personal Service, Food Production, English Language, Transportation, Public Safety and Security Oral Comprehension, Oral Expression, Near Vision, Problem Sensitivity, Speech Clarity
Engineer Critical Thinking, Reading Comprehension, Science, Active Listening, Complex Problem Solving Engineering and Technology, Mathematics, Design, Physics, Computers and Electronics Written Comprehension, Deductive Reasoning, Inductive Reasoning, Information Ordering, Problem Sensitivity
Executive Judgment and Decision Making, Complex Problem Solving, Critical Thinking, Coordination, Management of Financial Resources Administration and Management, Personnel and Human Resources, Customer and Personal Service, English Language, Economics and Accounting Oral Comprehension, Oral Expression, Speech Clarity, Written Comprehension, Deductive Reasoning
Film Director Active Listening, Critical Thinking, Monitoring, Reading Comprehension, Speaking Communications and Media, English Language, Telecommunications, Computers and Electronics, Administration and Management Oral Expression, Deductive Reasoning, Oral Comprehension, Problem Sensitivity, Speech Clarity
Firefighter Critical Thinking, Coordination, Judgment and Decision Making, Service Orientation, Active Learning Public Safety and Security, Customer and Personal Service, Education and Training, Building and Construction, English Language Problem Sensitivity, Oral Comprehension, Arm-Hand Steadiness, Deductive Reasoning, Far Vision
Guitar Player Speaking, Active Listening, Monitoring, Reading Comprehension, Social Perceptiveness Fine Arts, English Language, Foreign Language, Communications and Media, Education and Training Oral Comprehension, Oral Expression, Hearing Sensitivity, Auditory Attention, Memorization
Home Inspector Active Listening, Reading Comprehension, Speaking, Critical Thinking, Complex Problem Solving Building and Construction, Customer and Personal Service, Mathematics, Engineering and Technology, Design Problem Sensitivity, Inductive Reasoning, Deductive Reasoning, Oral Comprehension, Oral Expression
Hunter Critical Thinking, Operation and Control, Active Listening, Judgment and Decision Making, Operations Monitoring Law and Government, Mechanical, Geography, Production and Processing, Biology Arm-Hand Steadiness, Manual Dexterity, Multilimb Coordination, Static Strength, Extent Flexibility
Investigator Active Listening, Speaking, Critical Thinking, Complex Problem Solving, Reading Comprehension Law and Government, Public Safety and Security, English Language, Customer and Personal Service, Psychology Inductive Reasoning, Oral Comprehension, Deductive Reasoning, Oral Expression, Problem Sensitivity
Janitor Active Listening, Speaking, Coordination, Critical Thinking, Monitoring Public Safety and Security, Administration and Management, English Language, Customer and Personal Service, Transportation Near Vision, Trunk Strength, Arm-Hand Steadiness, Extent Flexibility, Manual Dexterity
Journal Editor Reading Comprehension, Writing, Active Listening, Critical Thinking, Speaking English Language, Communications and Media, Administration and Management, Administrative, Education and Training Written Comprehension, Written Expression, Oral Comprehension, Oral Expression, Fluency of Ideas
Journalist Active Listening, Reading Comprehension, Speaking, Writing, Critical Thinking English Language, Communications and Media, Law and Government, Computers and Electronics, Telecommunications Speech Clarity, Oral Expression, Oral Comprehension, Written Comprehension, Written Expression
Judge Active Listening, Critical Thinking, Judgment and Decision Making, Reading Comprehension, Complex Problem Solving Law and Government, English Language, Administration and Management, Psychology, Customer and Personal Service Deductive Reasoning, Oral Comprehension, Written Comprehension, Inductive Reasoning, Oral Expression
Lawyer Active Listening, Speaking, Reading Comprehension, Critical Thinking, Complex Problem Solving Law and Government, English Language, Customer and Personal Service, Administration and Management, Personnel and Human Resources Oral Expression, Oral Comprehension, Written Comprehension, Speech Clarity, Written Expression
Lifeguard Monitoring, Speaking, Social Perceptiveness, Service Orientation, Active Listening Customer and Personal Service, Public Safety and Security, English Language, Education and Training, Medicine and Dentistry Problem Sensitivity, Far Vision, Oral Expression, Oral Comprehension, Selective Attention
Manager Speaking, Reading Comprehension, Active Listening, Coordination, Writing Customer and Personal Service, Administration and Management, Economics and Accounting, English Language, Law and Government Oral Comprehension, Oral Expression, Written Comprehension, Written Expression, Inductive Reasoning
Mechanic Active Listening, Critical Thinking, Reading Comprehension, Complex Problem Solving, Speaking Engineering and Technology, Mechanical, Design, Mathematics, English Language Oral Comprehension, Written Comprehension, Information Ordering, Near Vision, Deductive Reasoning
Model Social Perceptiveness, Active Listening, Speaking, Coordination, Critical Thinking Customer and Personal Service, English Language, Fine Arts, Transportation, Communications and Media Oral Comprehension, Gross Body Coordination, Gross Body Equilibrium, Oral Expression, Speech Clarity
Nurse Social Perceptiveness, Active Listening, Coordination, Critical Thinking, Service Orientation Psychology, Customer and Personal Service, Medicine and Dentistry, English Language, Administrative Deductive Reasoning, Problem Sensitivity, Inductive Reasoning, Oral Comprehension, Oral Expression
Photographer Active Listening, Speaking, Service Orientation, Active Learning, Complex Problem Solving Customer and Personal Service, Sales and Marketing, Computers and Electronics, Administration and Management, Communications and Media Near Vision, Far Vision, Oral Expression, Originality, Visualization
Piano Player Speaking, Active Listening, Monitoring, Reading Comprehension, Social Perceptiveness Fine Arts, English Language, Foreign Language, Communications and Media, Education and Training Oral Comprehension, Oral Expression, Hearing Sensitivity, Auditory Attention, Memorization
Pilot Operation and Control, Operations Monitoring, Critical Thinking, Monitoring, Active Listening Transportation, Customer and Personal Service, Geography, English Language, Public Safety and Security Control Precision, Far Vision, Near Vision, Problem Sensitivity, Response Orientation
Plumber Critical Thinking, Judgment and Decision Making, Repairing, Troubleshooting, Monitoring Building and Construction, Mechanical, Design, Mathematics, Customer and Personal Service Problem Sensitivity, Finger Dexterity, Near Vision, Deductive Reasoning, Manual Dexterity
Poet Writing, Reading Comprehension, Active Listening, Critical Thinking, Speaking English Language, Communications and Media, Psychology, Administrative, Sales and Marketing Written Expression, Fluency of Ideas, Originality, Written Comprehension, Near Vision
Politician
Professor Reading Comprehension, Instructing, Writing, Speaking, Active Listening English Language, Education and Training, Communications and Media, History and Archeology, Philosophy and Theology Written Comprehension, Oral Expression, Oral Comprehension, Written Expression, Speech Clarity
Programmer Programming, Active Listening, Complex Problem Solving, Critical Thinking, Quality Control Analysis Computers and Electronics, Mathematics, Engineering and Technology, English Language, Customer and Personal Service Written Comprehension, Near Vision, Oral Comprehension, Deductive Reasoning, Inductive Reasoning
Research Assistant Critical Thinking, Active Listening, Reading Comprehension, Speaking, Writing English Language, Mathematics, Customer and Personal Service, Administration and Management, Computers and Electronics Inductive Reasoning, Written Comprehension, Written Expression, Deductive Reasoning, Mathematical Reasoning
Researcher Critical Thinking, Active Listening, Reading Comprehension, Speaking, Writing English Language, Mathematics, Customer and Personal Service, Administration and Management, Computers and Electronics Inductive Reasoning, Written Comprehension, Written Expression, Deductive Reasoning, Mathematical Reasoning
Salesperson Active Listening, Speaking, Negotiation, Persuasion, Social Perceptiveness Sales and Marketing, Customer and Personal Service, English Language, Mathematics, Transportation Oral Expression, Oral Comprehension, Speech Clarity, Speech Recognition, Written Comprehension
Scientist Complex Problem Solving, Critical Thinking, Judgment and Decision Making, Active Listening, Reading Comprehension Computers and Electronics, Mathematics, Engineering and Technology, English Language, Administration and Management Deductive Reasoning, Inductive Reasoning, Oral Comprehension, Oral Expression, Fluency of Ideas
Secretary Active Listening, Speaking, Reading Comprehension, Writing, Service Orientation Administrative, English Language, Computers and Electronics, Customer and Personal Service, Administration and Management Oral Comprehension, Oral Expression, Written Comprehension, Written Expression, Near Vision
Senator
Singer Speaking, Active Listening, Monitoring, Reading Comprehension, Social Perceptiveness Fine Arts, English Language, Foreign Language, Communications and Media, Education and Training Oral Comprehension, Oral Expression, Hearing Sensitivity, Auditory Attention, Memorization
Supervisor Active Listening, Management of Personnel Resources, Monitoring, Speaking, Coordination Customer and Personal Service, Administration and Management, English Language, Personnel and Human Resources, Economics and Accounting Oral Comprehension, Oral Expression, Speech Recognition, Speech Clarity, Deductive Reasoning
Surgeon Complex Problem Solving, Judgment and Decision Making, Critical Thinking, Reading Comprehension, Active Learning Medicine and Dentistry, Biology, English Language, Customer and Personal Service, Psychology Arm-Hand Steadiness, Finger Dexterity, Near Vision, Control Precision, Deductive Reasoning
Tailor Time Management, Active Listening, Critical Thinking, Speaking, Social Perceptiveness Customer and Personal Service, English Language, Production and Processing, Administration and Management, Economics and Accounting Arm-Hand Steadiness, Finger Dexterity, Visualization, Near Vision, Oral Comprehension
Teacher Instructing, Speaking, Learning Strategies, Active Listening, Critical Thinking Education and Training, English Language, Mathematics, Psychology, Computers and Electronics Oral Expression, Deductive Reasoning, Oral Comprehension, Problem Sensitivity, Speech Clarity
Technician Critical Thinking, Reading Comprehension, Complex Problem Solving, Active Listening, Troubleshooting Computers and Electronics, Engineering and Technology, English Language, Design, Mathematics Problem Sensitivity, Deductive Reasoning, Near Vision, Inductive Reasoning, Written Comprehension
Violin Player Speaking, Active Listening, Monitoring, Reading Comprehension, Social Perceptiveness Fine Arts, English Language, Foreign Language, Communications and Media, Education and Training Oral Comprehension, Oral Expression, Hearing Sensitivity, Auditory Attention, Memorization
Writer Writing, Reading Comprehension, Active Listening, Speaking, Critical Thinking Sales and Marketing, Communications and Media, Customer and Personal Service, Computers and Electronics, Mathematics Written Expression, Written Comprehension, Oral Comprehension, Oral Expression, Fluency of Ideas